Problem (4): Simplifying boolean formulae (30 pts) uestion (a): Two short ones (5 pts each for 10 pts total) Using the axioms of propositional logie equivalence outlined in lecture and available in your book, simplify the following compound statements as much as possible. For every derivation you make, write down the axiom you apply on the right of the derivation (e.g law of associativity, De Morgans, ete). BEGIN YOUR ANSWER TO QUESTION (a) BELOW THIS LINE
Answer #1

(i)

(k ^l ) v ((k ^ m) ^ k ^ (l v ~m))

(k ^ l) v ((k ^ k) ^ m ^ (l v ~m)) Associative law ((A^B) ^ C = (A^C) ^ B)

(k ^l ) v (k ^ m ^ (l v ~m)) Idempotent law (k ^ k =k)

(k ^ l) v (k ^ (m ^ l) v (m ^ ~m)) Distibutive law   ((A^B) v C = (A^C) v ( B ^C))

(k ^ l) v (k ^ (m ^ l) v 0) Compliment law ( A ^ ~A = 0)

(k ^ l) v (k ^ (m ^ l)) Identity law (A v 0 = A)

(k ^ l) v ((k ^ l) ^ m) Associative law ((A^B) ^ C = (A^C) ^ B)

(k ^ l) Absorption law (A v (A ^ B) = A)

(ii)

z ^ (q v (~w ^ q )) ^ (~z v ~q)

z ^ q ^ (~z v ~q) Absorption law (A v (A ^ B) = A)

(z ^ q) ^ ~(z ^ q) Demogan Law (~(A ^ B) = (~A v ~B ) )

0 Compliment law ( A ^ ~A = 0)